# Enhanced SPS Velocity-Adaptive Scheme: Access Fairness in 5G NR V2I Networks

### v1

@@ 9       Selection: Select the top $M$ individuals in $Q_{n}$ as $P_{n+1}$ using $\bm{F}$ and $d(\bm{w_{i}})$;
11Optimization Phase: ;
12 $P_{\text{filtered}}=\bm{w}\in$$P_{n+1}$ : $F_{K_{i}}(\bm{w})\leq\text{Threshold},\forall i$;
13 $\bm{w}^{*}=\arg\min_{\bm{w}\in P_{\text{filtered}}}\sum_{i=1}^{N}F_{K_{i}}(\bm{w})$;
14
return $\bm{w}^{*}$
@@ 13 $\bm{w}^{*}=\arg\min_{\bm{w}\in P_{\text{filtered}}}\sum_{i=1}^{N}F_{K_{i}}(\bm{w})$;
We use the NSGA-II algorithm [17] to address the optimization problem. Each population consists of $N$ individuals, representing the vehicle selection windows. The population size is $M$. The algorithm can be organized into three parts: initialization phase, iteration phase, and optimization phase. The algorithm’s procedure is presented in Algorithm 1.
2) Iteration Phase: Firstly, crossover and mutation are performed. Crossover involves exchanging parts of two parent individuals’ genes to generate new offspring, while mutation randomly modifies an individual’s genes to prevent the algorithm from converging to a local optimum. At this point, the parent and offspring populations are merged into a new population $Q_{n}$ for the subsequent selection operation. Next, we use $\bm{v}=\{v_{1},v_{2},\dots,v_{N}\}$ to calculate the objective function values $F_{K}(\bm{w})$ according to Eq.(21). Then we perform non-dominated sorting. Specifically, if individual $A$ is equal to or better than individual $B$ in every objective, and better than $B$ no fewer than one objective, individual $A$ is regarded as dominate individual $B$. Non-dominated sorting divides the population into multiple levels according to each individual’s dominance rank. Thus we derive $\bm{F}=[F_{1},F_{2},\dots]$, $F_{1}$ means the individuals which rank first, and then, crowding distance is calculated as follow[37]:
$d(\bm{w_{i}})=\sum_{j=1}^{N}\left(\frac{f_{j}^{i+1}(\bm{w_{i}})-f_{j}^{i-1}(\bm{w_{i}})}{f_{j}^{max}-f_{j}^{min}}\right),$ (22)
where $f_{j}^{i+1}$ and $f_{j}^{i-1}$ respectively represent the objective value on objective $j$ for the previous and next individual of the dominance rank individual $i$. $f_{j}^{max}$ and $f_{j}^{min}$ respectively represents the maximum and minimum objective values of all individuals on the objective $j$[38].
Then, parent individuals are selected based on non-dominated sorting and crowding distance. The selection prefers individuals that are ranked higher in non-dominated sorting, and within the same rank, individuals with larger crowding distances are preferred. Thus, the merged population $Q_{n}$ undergoes non-dominated sorting and selection. Finally, $M$ individuals with higher domination rank and larger crowding distances are selected to form the new population $p_{n+1}$.
3) Optimization Phase: In this phase, the optimal population is selected from the Pareto front. The optimal population must ensure that the difference between the fairness index of each vehicle is smaller than a threshold, while also minimizing the total objective function values.
Thus, the optimal selection window $\bm{w}^{*}$ is obtained.
@@ VI Numerical Simulation And Analysis
In this section, we validate the effectiveness of our proposed scheme. The simulation experiments are implemented using Python 3.9. The scenario is set as a two-way highway with four lanes, where the speed limit is between 20 and 30 m/s, and the speed difference between adjacent lanes is within 4 m/s. The window size limits and time slot length are configured according to the 5G NR V2I Mode 2 settings.
@@
Figure 5: Optimal Selection Window Versus Average Velocity
Fig. 3 presents the performance metrics of the NSGA-II for solving the proposed multi-objective problem, evaluated using hypervolume (HV), Inverted Generational Distance (IGD), Generational Distance (GD), and Spacing.
The hypervolume (HV) measures the extent to which the solution set covers the objective space, progressively increasing and stabilizing, indicating high diversity. Generational Distance (GD) evaluates the proximity of the obtained solutions to the true Pareto front, with values steadily decreasing and stabilizing, demonstrating effective convergence. Inverted Generational Distance (IGD) considers both diversity and convergence, yielding higher values than GD. Lastly, Spacing measures the uniformity of solution distribution, remaining consistently low and indicating an even spread of solutions.
@@
In Fig. 5, we present a comparison of the objective values when vehicles adopt the optimal selection window versus the standard selection window. we can see that when using the standard selection window, the objective values of vehicles progressively increase, indicating a gradual loss of fairness in access. This phenomenon arises because faster vehicles will have less staying time in RSU coverage, resulting in transmitting less data, leading to unfairness. However, when the optimal selection window is employed, both the rate of increase and the initial values are significantly smaller compared to the standard window. This improvement is attributed to the vehicles’ ability to adaptively adjust the optimal window based on their speed, thereby striving to maintain fairness.

### v3

@@ 7       Non-dominated Sorting: Compute $\boldsymbol{F}=\{F_{1},F_{2},\dots\}$ for $Q_{n}$;
[…]
12 $\boldsymbol{w}^{*}=\arg\min_{\boldsymbol{w}\in P_{\text{filtered}}}\sum_{i=1}^{N}F_{K_{i}}(\boldsymbol{w})$;
@@ VI Numerical Simulation And Analysis
In this section, we validate the effectiveness of our proposed scheme.The experimental setup and implementation details can be accessed from the source code repository at https://github.com/qiongwu86/Enhanced-SPS-Velocity-adaptive-Scheme-Access-Fairness-in-5G-NR-V2I-Networks. The scenario is set as a two-way highway with four lanes, where the speed limit is between 20 and 30 m/s, and the speed difference between adjacent lanes is within 4 m/s. The window size limits and time slot length are configured according to the 5G NR V2I Mode 2 settings.
Fig. 3 presents the performance metrics of the NSGA-II using hypervolume (HV), Inverted Generational Distance (IGD), Generational Distance (GD), and Spacing. The HV measures the extent to which the solution set covers the objective space, progressively increasing and stabilizing, indicating high diversity. GD evaluates the proximity of the obtained solutions to the true Pareto front, with values steadily decreasing and stabilizing, demonstrating effective convergence. IGD considers both diversity and convergence, yielding higher values than GD. Lastly, Spacing measures the uniformity of solution distribution, remaining consistently low and indicating an even spread of solutions.
@@
Figure 5: Objective Function Value Versus Average Velocity
@@
In Fig. 5, we present a comparison of the objective values when vehicles adopt the optimal selection window versus the standard selection window. It can be observed that when using the standard selection window, the objective values of vehicles progressively increase, indicating a gradual loss of fairness in access. This phenomenon arises because faster vehicles will have less staying time in RSU coverage, resulting in transmitting less data, leading to unfairness. However, when the optimal selection window is employed, both the rate of increase and the initial values are significantly smaller compared to the standard window. This improvement is attributed to the vehicles’ ability to adaptively adjust the optimal window based on their speed, thereby striving to maintain fairness.
@@ VII Conclusion
Acknowledgment
This work was supported in part by Jiangxi Province Science and Technology Development Programme under Grant No. 20242BCC32016, in part by the National Natural Science Foundation of China under Grant No. 61701197, in part by the National Key Researchand Development Program of China under Grant No. 2021YFA1000500(4) and in part by the 111 Project under Grant No. B23008. (Corresponding author: Qiong Wu.)